• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name fissum

    Log entry time 21:19:41 on March13,2001

    Entry number 58714

    keyword=target instability due to Hall C cooldown

    21.05 Asked to reduce the beam current due to the Hall C
    target cooldown removing our overhead. Started reducing
    the beam in steps of 10 uA, but then decided to turn it off.
    Came back at 40 uA. Still no good. Kathy suggested to talk
    to Hall C and see if they might be ready for the tuning now
    instead of at 22.00. That way tuning and cooldown (no beam
    to us) would all occur at the same time, and we could get
    running sooner. Stay tuned...